# Uncertainty Quantification and Confidence Calibration in Large Language Models: A Survey

### v1

@@ 6. Challenges and Future Directions
While significant strides have been made in integrating uncertainty quantification into Large Language Models, several unaddressed challenges persist. This section will explore these unresolved issues, ranging from efficiency-performance trade-offs to cross-modal uncertainty, and outline promising avenues for future research, aiming to advance the reliability and trustworthiness of LLMs in high-stakes applications.
@@
Efficiency-Performance Trade-offs. Multi-sample uncertainty methods incur prohibitive costs for trillion-parameter LLMs (\$12k per million queries (Li et al., 2024a)), yet yield marginal reliability gains ($\leq 0.02$ AUROC improvement (Xiong et al., 2024b)). Hybrid approaches combining low-cost proxies (attention variance (Heo et al., 2018), hidden state clustering (Nikitin et al., 2024)) could resolve this by achieving 90% of maximal performance at 10% computational cost. For example, precomputing uncertainty “hotspots” during inference could trigger targeted multi-sampling only for high-risk outputs like medical diagnoses.
@@
Interpretability Deficits. Users struggle to distinguish whether uncertainty stems from ambiguous inputs (“Explain quantum gravity”), knowledge gaps, or decoding stochasticity—a critical barrier in high-stakes domains. Modular architectures that decouple uncertainty estimation layers (Huang et al., 2025b; Sensoy et al., 2018) or employ causal tracing of transformer attention pathways (Wang et al., 2024b) could clarify uncertainty origins. For instance, perturbing model weights (Gal and Ghahramani, 2016) might reveal parametric uncertainty in low-resource languages, while input modules flag underspecified terms (”this country”) for clarification.
@@
Cross-Modality Uncertainty. Integrating vision, text, and sensor data introduces misaligned confidence estimates between modalities: LVLMs exhibit 2.4×2.4$\times$2.4 × higher uncertainty in visual vs. textual components (Zhang et al., 2024b), causing 63% of errors in multi-modal QA (Zhang et al., 2024d). Dynamic contrastive decoding and uncertainty-aware fusion protocols show promise(Huo et al., 2024; Suo et al., 2025), but require domain-specific adaptations (e.g., aligning radiology images with reports (Li et al., 2023b; Zhao et al., 2023b)). Future work must develop unified uncertainty embeddings to harmonize modality confidence scales and adversarial training against cross-modal backdoor attacks (Zhang et al., 2024c; Walmer et al., 2022; Liang et al., 2024a).
Interventions for Uncertainty. Real-time uncertainty monitoring enables self-correction (e.g., retracting erroneous reasoning steps (Yin et al., 2024)), but adds 300 to 500ms latency per decoding iteration (Xiong et al., 2024b). Lightweight monitors via attention variance distillation (Gong et al., 2021; Huang et al., 2022) and human-in-the-loop calibration (Wang et al., 2022) could mitigate overhead while addressing security risks like adversarial uncertainty manipulation (100% attack success rates (Zeng et al., 2024)). Certified defense mechanisms and causal propagation tracing (Goyal et al., 2023) are critical for deploying uncertainty-guided interventions in sensitive applications.
UQ Evaluation. There are several challenges to effectively evaluating UQ quality. There are limitations in adopting a binary classification evaluation approach: It is challenging to decide if a free-form generation is correct even in a QA dataset, let alone more open-ended tasks. Also, even LLM-as-a-judge suffers systematic biases including self-preference bias (where an LLM judge may favor answers written in a style similar to its own) (Panickssery et al., 2024), length bias (where LLM judges often prefer longer, more detailed answers) (Lin et al., 2022b), and position bias (where the order of presented answers influences the judgment) (Zheng et al., 2023b). Moreover, current evaluation methods, such as AUROC and AUARC, fall short in capturing nuanced, “meaningful” uncertainty. For instance, they often do not differentiate between cases where a model is confidently wrong (hallucinations) and cases where it expresses a prudent level of uncertainty. Finally, a significant practical challenge in evaluating UQ for LLMs is the lack of suitable datasets that align with what we want to measure. Many benchmarks exist to test LLM capabilities (from reading comprehension to factual QA), but few are explicitly designed for UQ.
@@ 7. Conclusion
In this survey, we offer a comprehensive overview of uncertainty quantification (UQ) in Large Language Models (LLMs). Initially, we introduce the fundamental concepts relevant to both UQ and LLMs, highlighting the importance of reliability in high-stakes applications. Following this, we propose a detailed taxonomy for characterizing uncertainty dimensions in LLMs, including input, reasoning, parameter, and prediction uncertainty. In terms of methodologies, we systematically evaluate UQ methods using our novel taxonomy, thoroughly reviewing their effectiveness across different uncertainty types. Ultimately, we identify and discuss some of the persistent challenges in UQ for LLMs, providing insightful directions for future research. The primary goal of this survey is to promote a more seamless integration of UQ techniques into LLM development, motivating both machine learning researchers and practitioners to delve into this rapidly advancing area.

### v2

@@ 6. Challenges and Future Directions
While significant strides have been made in integrating uncertainty quantification into LLMs, several unaddressed challenges persist. This section will explore these unresolved issues, ranging from efficiency-performance trade-offs to cross-modal uncertainty, and outline promising avenues for future research, aiming to advance the reliability of LLMs in high-stakes applications.
@@
$\bullet$ Efficiency-Performance Trade-offs. Multi-sample uncertainty methods incur prohibitive costs for trillion-parameter LLMs (\$12k per million queries (Li et al., 2024a)), yet yield marginal reliability gains ($\leq 0.02$ AUROC improvement (Xiong et al., 2024b)). Hybrid approaches combining low-cost proxies (attention variance (Heo et al., 2018), hidden state clustering (Nikitin et al., 2024)) could resolve this by achieving 90% of maximal performance at 10% computational cost. For example, precomputing uncertainty “hotspots” during inference could trigger targeted multi-sampling only for high-risk outputs like medical diagnoses.
@@
$\bullet$ Interpretability Deficits. Users struggle to distinguish whether uncertainty stems from ambiguous inputs, knowledge gaps, or decoding stochasticity. Modular architectures that decouple uncertainty estimation layers (Huang et al., 2025b; Sensoy et al., 2018) or employ causal tracing of transformer attention pathways (Wang et al., 2024b) could clarify uncertainty origins. For instance, perturbing model weights (Gal and Ghahramani, 2016) might reveal parametric uncertainty in low-resource languages, while input modules flag underspecified terms for clarification.
@@
$\bullet$ Cross-Modality Uncertainty. Integrating vision, text, and sensor data introduces misaligned confidence estimates between modalities: LVLMs exhibit 2.4×2.4$\times$2.4 × higher uncertainty in visual vs. textual components (Zhang et al., 2024a), causing 63% of errors in multi-modal QA (Zhang et al., 2024d). Dynamic contrastive decoding and uncertainty-aware fusion protocols show promise(Huo et al., 2024; Suo et al., 2025), but require domain-specific adaptations (e.g., aligning radiology images with reports (Li et al., 2023b; Zhao et al., 2023b)). Future work must develop unified uncertainty embeddings to harmonize modality confidence scales and adversarial training against cross-modal backdoor attacks (Zhang et al., 2024c; Liang et al., 2024a).
$\bullet$ System-level Uncertainty in Agents and Reasoning. As LLMs are increasingly deployed as autonomous agents or reasoning engines, the propagation and accumulation of uncertainty across steps becomes critical. Errors in early steps can lead to cascading failures, especially when the model expresses misplaced confidence. However, most existing UQ methods operate at one round of outputs from LLM, lacking mechanisms to capture uncertainty over multi-step reasoning chains or multi-action plans. As studies suggest that LLMs often fail to revise earlier decisions when presented with contradictory information (Creswell et al., [n. d.]), there is a need for temporally-aware uncertainty tracking. Enhancing LLMs with structured memory or model-based planning, or leveraging graph-based representations to trace and revise uncertain steps (Madaan et al., 2023; Yao et al., 2023), could possibly provide more reliable behavior.
$\bullet$ UQ Evaluation. Evaluating the quality of UQ remains a fundamental challenge. While the binary classification metrics introduced in Section 4.2 are widely used, they are not always suitable: many tasks, especially in NLG, cannot be easily reduced to binary correctness. Even for structured tasks like question answering, determining whether a free-form generation is correct can be nontrivial due to semantic variability and ambiguity. This issue becomes even more pronounced in open-ended tasks. Moreover, LLM-as-a-judge evaluation approaches are themselves subject to systematic biases (Panickssery et al., 2024; Lin et al., 2022b; Zheng et al., 2023). In addition, common evaluation metrics such as AUROC and AUARC often fail to capture what might be considered “meaningful” uncertainty. These metrics typically assess a model’s ability to distinguish between correct and incorrect outputs, but do not differentiate between confidently wrong responses and those accompanied by an appropriate level of uncertainty.
@@ 7. Conclusion
In this survey, we offer a comprehensive overview of uncertainty quantification (UQ) in Large Language Models (LLMs). We first introduce the fundamental concepts relevant to both UQ and LLMs, highlighting the importance of reliability in high-stakes applications. Following this, we propose a detailed taxonomy for characterizing uncertainty dimensions in LLMs, including input, reasoning, parameter, and prediction uncertainty. We systematically introduce existing UQ methods using our novel taxonomy, reviewing their effectiveness across different uncertainty types. Ultimately, we identify and discuss some of the persistent challenges in UQ for LLMs, providing insightful directions for future research. The primary goal of this survey is to promote the integration of UQ techniques into LLM development, motivating both machine learning researchers and practitioners to participate in this rapidly advancing area.
ACKNOWLEDGMENTS
The work was partially supported by NSF awards #2421839. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies. We thank Amazon Research Awards and OpenAI for providing us with API credits under the Researcher Access program.